OpenStudy (anonymous):

If a vertex is (3,0) and focus at (3,-2), what is the equation of the parabola?

OpenStudy (anonymous):

is it \[(x-3)^{2} =8y\] or \[(x-3)^{2} =-8y\]

OpenStudy (tkhunny):

Well, if it is "+", how is it going to go down around that focus?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ \text{Okay if you don't believe me Let's work backwards }\\ (x-3)^2=8y\\ y= \frac{(x-3)^2}{8}\\ y=a(x-h)^2+k \ \ \ Focus=(3,2) \ \ \ \ Directrix: y=-2 , \\ Vertex: (3,0), \ \ Axis\ \ \ of\ \ Symmetry: x=3\]

OpenStudy (anonymous):

Got it Thanks! I had gotten positive, but my friend's note said negative, So i was a bit confused.. and then I thought that the parabola had to be graphed upwards in order to overlap with the focus.. Since the graph looks like that, I thought it had to include the focus point of (3,-2)

OpenStudy (anonymous):

you did well. :)